OpenStudy (anonymous):

Solve for x. 8/x-7 = 7/2

OpenStudy (anonymous):

@TheDrowzeeAbra

OpenStudy (anonymous):

x=16/21

OpenStudy (anonymous):

add seven to both sides 8/x = 7+7/2 =?

OpenStudy (anonymous):

A. 9 2/7 B. x = 31 1/2 C. x = 35 D. x = 63

OpenStudy (anonymous):

it had to be one of those

OpenStudy (anonymous):

@hk32176